The driver of a train running at 25m/s sights a train B moving in the same direction on the same track with 15m/s.Thedriver of the train ‘A’ applies breaks to produce decelaration of 1m/s2. what should be the minimum distace between the trains to avoid accident?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From first eqn.
= 25 – 1*t
t = 25 seconds. the Train A would stop.
and distance it would cover in that time,
0^2 = 625  – 2*1*s
s = 625/2 m = > 312.5 m.
in that time, train would have covered,
x  = 15*25 = > 375 meter.
The minimum distance should be=> 375 – 312.5 = 62.5 m.